# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v2

@@ 4.2 Infrared
We detected AT2019aalc in all of the three bands observed with the WIRC. The observed magnitudes are plotted in the top panel of Fig. 3. We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (Cutri et al. 2003). The observed magnitudes in April are higher with mag in J-, and Ks-bands and with mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
The top panel of Fig. 3 shows the light curve starting just before the first optical flare. The distinct first dust echo flare that peaks around six months after the first optical peak. The rise of the dust echo of the second flare becomes also visible in the latest WISE data releases. The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (Berk et al. 2004). In contrast, the dust echo presents an increase of mag in W1 and mag in W2 after the first optical flare. Following the second optical flare, the IR light curve shows a brightening of around mag in both bands and is increasing when the optical flare is already decaying. Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (Necker et al. 2024), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with . Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (Stern et al. 2012). However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (Stern et al. 2012), the color of AT2019aalc suggests a reliability of only %. With the onset of the dust echo, the IR emission becomes dominated by the heated dust. The change from bluer to redder colors comes from cooling of the dust. The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue shown in Fig.5.
Notably, AT2019aalc is not part of the Flaires sample (Necker et al. 2024), a list of dust-echo-like infrared flares interpreted as extreme AGN accretion events or TDEs. This is because the broad application of the Flaires pipeline necessarily included a strict cut on extraneous variability that AT2019aalc did not pass. The authors did note however that a desirable relaxation of this criterion would include AT2019aalc. The source is, however, part of the dust echo sample of van Velzen et al. (2024) and, notably has the highest dust echo flux of all ZTF transients.
@@ 4.3 X-ray
The X-ray luminosity evolution is shown in the bottom panel of Fig. 3 while the count rate variability is plotted in the top panel of Fig. 6. The first flaring episode peaked at in the end of June (see Fig. 3.) This flare peaked weeks before the optical peak. The count rate dropped by a factor of within a timescale of days. Later, the X-ray emission shows limited variability for several months. A hint of increased emission was detected in mid-February which was also followed by a bump in the optical and UV light curves around weeks later. In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s which is almost a magnitude larger compared to the peak of the first X-ray flaring episode. Interestingly, this X-ray flare was also accompanied by a bump in the optical weeks later. A similarly luminous and rapid flare was detected in August and once more, accompanied by an optical bump weeks later. Around days passed between the second and third and the third and fourth X-ray peaks.
@@
AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (Sunyaev et al. 2021) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence. The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV. The source had a soft thermal spectrum described with a blackbody temperature of eV (van Velzen et al. 2024). One observation with XMM-Newton Observatory (Jansen et al. 2001) was performed on 2021-02-21 during a period of optical quiescence. The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (Pasham 2023).
@@ 4.4.1 Temporal evolution
The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, White et al. 1997) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability. These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz which is typical for Seyfert galaxies (Meurs & Wilson 1984).
Later, the transient’s sky location was observed in the Karl G. Jansky Very Large Array Sky Survey (VLASS, Lacy et al. 2020). The VLASS observes the entire sky visible to the VLA at GHz wide band three times. Our source was detected with peak flux density values of mJy/beam (on 2019-03-14), mJy/beam (on 2021-11-06) and mJy/beam (on 2024-07-20), corresponding to 3 GHz radio luminosities () of , and , respectively. This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, van Velzen et al. 2024).
In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: McConnell et al. 2020). The source was observed in two different fields and, consequently, at two different epochs. The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare seen in the VLASS data started at least days after the peak of the first optical flare, sometime between April 2020 and November 2021.
The radio luminosity in the beginning of our ATCA monitoring was at the same level ( ) when the source was observed for the second time in the VLASS in April 2021. Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increasing resulted in a long-lasting plateau. We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig. 3.
@@ 4.4.2 The radio spectrum
To further characterize the radio emission of AT2019aalc, we derived the radio spectral index of the radio source in each epoch of our ATCA monitoring. The spectral index between GHz and GHz (, defined as ) does not show any obvious signs of variability on a timescale of months. We give an average spectral index of . We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities . The spectral index we derived between GHz and GHz this way is .
@@ 4.4.3 VLBI detection
The naturally-weighted VLBI map of AT2019aalc is presented in Fig. 9. We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of . We obtained the following parameters of the fitted component: a total flux density of mJy and a size of mas at a position angle of . We calculated the uncertainties of the flux density and size parameters following the formulas of Fomalont (1999). The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following Kovalev et al. (2005). This yields mas implying that the fitted Gaussian component is resolved.
We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Condon et al. 1982; Ulvestad et al. 2005):
where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz. The obtained brightness temperature is K at GHz. The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., Lobanov 2015).
@@ 4.5 Spectroscopic results
We present the archival and the newly obtained optical spectra of AT2019aalc in Fig. 17 with marking the notable emission lines. In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and five spectra that were taken during the second flare (over-flare-2 spectra).
@@ 4.5.1 Pre-flare spectrum
A spectrum of the pre-flare system was taken in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
We identify strong Balmer emission lines. We calculated FWHM(Balmer) kms-1 for the broad component of the Balmer lines in the pre-flare spectrum, which is typical for an unobscured, broad-line AGN. The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (Liu et al. 2019) for our source are consistent with this classification.
Several forbidden lines such as [O iii] and [O iii] , [S ii] , [S ii] and [N ii] were detected. These lines are produced in the narrow-line region (NLR) in AGN. We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., Santoro et al. 2020).
The pre-flare spectrum also shows the detections of two high-ionization coronal lines ([Fe vii] and [Fe xiv]) which are further discussed below.
@@ 4.5.2 Post-flare-1 spectrum
The first LRIS spectrum was taken almost years after the first optical flare and years before the second one.
@@
We identify the Bowen Fluorescence line O iii and He ii .
The high ionization coronal lines [Fe vii], [Fe x], [Fe xiv] and [Fe vii] were detected. We identify the [Ne V] emission lines at and in the spectrum. Note that the pre-flare SDSS spectrum does not cover these wavelengths, therefore these lines are not necessarily related to the optical flares discussed here. These lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks and are sometimes considered as coronal lines (e.g., Cann et al. 2020).
@@ 4.5.3 Over-flare-2 spectra
We performed optical spectroscopic observations of AT2019aalc during the second flare. These are characterized by a blue continuum with several emission lines appearing in comparison to the spectra discussed above. We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in Pol & Wadadekar (2017) ,in Fig. 10. This comparison shows similarities but also significant differences discussed below.
Over the second optical flare, strong Balmer lines were detected. Two lines of the series (H and H) were detected only in the over-flare-2 spectra.
Over the second optical flare of AT2019aalc, we detected the Bowen Fluorescence line complex N iii +He ii . This complex was not detected in the earlier spectra. The relative strength of the BF complex to H reached a maximum on 2023-07-11, very close to the optical continuum peak (2023-07-13). As we detected the Bowen line N iii even days after the optical peak of the second flare, we conclude that it persists over long timescales. We also find hints of O iii which is also known as a Bowen line (Selvelli et al. 2007), however, due to a possible blending from Fe vii, its detection is questionable. Nevertheless, this Bowen line is created through the same channel (O1) as the above discussed O iii which supports the hint of detection. O iii was detected in some metal-rich TDEs before (Leloudas et al. 2019) such as AT2019dsg (Cannizzaro et al. 2021).
We detect the high-ionization coronal lines seen in the post-flare-1 LRIS spectrum over the second flare in each spectrum. Furthermore, a new coronal line appears: (Fe xi (ionization potential: eV). We also detected the [Ne V] emission lines at and in the second LRIS spectrum, as well as in the post-flare-1 spectrum.
@@ 5.1.1 Balmer lines
AT2019aalc exhibits prominent Balmer lines in each spectrum. In addition, a more complex structure of these lines appears in the late-time over-flare-2 spectrum. The Balmer series lines here exhibit a redward wing which, however, in some cases, might result from blending with other lines, such as H and the BF line N,iii . Interestingly, the post-flare-I spectrum shows similar structure of the Balmer line series.
@@ 5.1.2 Bowen fluorescence lines
Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation. The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV) as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (Netzer et al. 1985; Netzer 1990). This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc. The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot et al. 2019) and AT2021loi and only in a few Seyfert galaxies before (e.g., Malkan 1986; Schachter et al. 1990). However, we note that due to the atmospheric cutoff at these lines are difficult to detect. This line is typically weaker in regular Seyfert galaxies than in the case of AT2019aalc (see Fig. 10). The BF line N iii is not generally seen in AGN (e.g., Vanden Berk et al. 2001), however, it has been observed in several optical TDEs (Charalampopoulos et al. 2022b) and other SMBH-related transients: BFFs (Trakhtenbrot et al. 2019; Makrygianni et al. 2023), the peculiar transient event AT2019avd (Malyali et al. 2021) and a few flaring NLSy1s (Frederick et al. 2021).
The detection of a BF line almost years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot et al. (2019). In the case of AT2017bgt, this line was still detected days after the discovery of the transient. The Bowen features in F01004-2237 were detected even years after the initial flare (Tadhunter et al. 2021). We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc to the spectra of BFFs, a TDE-Bowen, and a repeating partial TDE in Fig. 11 around the N iii He ii complex. We further compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig. 12 zooming into the O iii He ii region.
The He ii ionized helium line is presented in the spectrum taken between the optical flares.. This emission line is commonly present in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g. Wang & Kron 2020). This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (Pakull & Angebault 1986; Schaerer et al. 2019; Cannizzaro et al. 2021). In the spectra of AT2019aalc, He ii is detected together with strong N iii and O iii BF lines and He ii , indicating its relation to the Bowen Fluorescence mechanism. Furthermore, we calculated a maximum line intensity ratio of which significantly exceeds the typical ratios estimated for AGN (earlier Vanden Berk et al. (2001) estimated ratios of while Shirazi & Brinchmann (2012) gives ). Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot et al. 2019).

### v3

@@ 4.2 Infrared
We detected AT2019aalc in all of the three bands observed with the WIRC. The observed magnitudes are plotted in the top panel of Fig. 3. We compare these to pre-flare magnitudes presented in the 2MASS All-Sky Catalog of Point Sources (2003tmc..book.....C). The observed magnitudes in April are higher with mag in J-, and Ks-bands and mag in H-band indicating the clear brightening of the IR counterpart with respect to archival data.
The top panel of Fig. 3 shows the light curve starting just before the first optical flare. The distinct first dust echo flare peaks around six months after the first optical peak. The rise of the dust echo of the second flare also becomes visible in the latest WISE data releases. The light curve varies by around mag in W1 and mag in W2 before the dust echo flare, which is consistent with statistical AGN variability (berkEnsemblePhotometricVariability2004). In contrast, the dust echo presents an increase of mag in W1 and mag in W2 after the first optical flare. Following the second optical flare, the IR light curve shows a brightening of around mag in both bands and is increasing when the optical flare is already decaying. Because the brightness is significantly above the limiting magnitude where Eddington bias becomes an issue for stacked single-exposure photometry (2025A&A...695A.228N), we can estimate the pre-flare mid-IR color using the median of the light curve prior to the dust echo flare with . Although there is clear AGN activity prior to the dust echo flare, this is still below the widely used AGN identification cut of (2012ApJ...753...30S). However, according to the reliability of WISE selection as a function of a simple W1W2 color selection (2012ApJ...753...30S), the color of AT2019aalc suggests a reliability of only %. With the onset of the dust echo, the IR emission becomes dominated by the heated dust. The change from bluer to redder colors comes from cooling of the dust. The re-brightening of the second dust echo is consequently accompanied by another color change towards the blue, shown in Fig. 5.
Notably, AT2019aalc is not part of the Flaires sample (2025A&A...695A.228N), a list of dust-echo-like infrared flares interpreted as extreme AGN accretion events or TDEs. This is because the broad application of the Flaires pipeline necessarily included a strict cut on extraneous variability that AT2019aalc did not pass. The authors did note however that a desirable relaxation of this criterion would include AT2019aalc. The source is, however, part of the dust echo sample of Lancel and, notably has the highest dust echo flux of all ZTF transients.
@@ 4.3 X-ray
The keV X-ray luminosity evolution is shown in the bottom panel of Fig. 3 while the count rate variability is plotted in the top panel of Fig. 6. The first flaring episode peaked at a luminosity of in the end of June (see Fig. 3.) This flare peaked two weeks before the optical peak. The count rate dropped by a factor of within a timescale of days. Later, the X-ray emission showed limited variability for several months. A hint of increased emission was detected in mid-February , which was also followed by a bump in the optical and UV light curves around two weeks later. In the end of April , a rapid and extreme flaring episode started in X-rays that lasted less than a month but reached a luminosity of erg/s. which is almost a magnitude larger than the peak of the first X-ray flaring episode. Interestingly, this X-ray flare was also accompanied by a bump in the optical weeks later. A similarly luminous and rapid flare was detected in August and once more, accompanied by an optical bump weeks later. Around days passed between the second and third and the third and fourth X-ray peaks.
@@
AT2019aalc (= SRGe J152416.7045118) was observed by the eROSITA (2021A&A...656A.132S) soft X-ray telescope four times starting from 2020-02-02 (roughly half a year after the peak of the first optical flare) with a months cadence. The X-ray light curve reached a plateau between August and January with a flux of F erg cm-2s-1 in the energy range of keV. The source had a soft thermal spectrum described with a blackbody temperature of eV (Lancel). One observation with XMM-Newton Observatory (2001A&A...365L...1J) was performed on 2021-02-21 during a period of optical quiescence. The best-fit power-law index and the observed keV flux of this observation were and F erg cm-2s-1 (68 confidence), respectively (2023ATel16118....1P).
@@ 4.4.1 Temporal evolution
The pre-flare system was detected in the NRAO VLA Sky Survey (NVSS, 1997ApJ...475..479W) in February 1995 and in the FIRST survey in February 2000 (both at GHz) with peak flux densities of mJy and mJy respectively, suggesting no substantial variability. These flux densities correspond to a monochromatic radio luminosity of W Hz-1 measured at GHz, which is typical for Seyfert galaxies (1984A&A...136..206M).
Later, the transient’s sky location was observed in the Karl G. Jansky Very Large Array Sky Survey (VLASS, 2020PASP..132c5001L). The VLASS observes the entire sky visible to the VLA at GHz wide band three times. Our source was detected with peak flux density values of mJy/beam (on 2019-03-14), mJy/beam (on 2021-11-06) and mJy/beam (on 2024-07-20), corresponding to 3 GHz radio luminosities () of , and , respectively. This factor of increase in flux from the first VLASS epoch (three months before the first optical peak) to the second VLASS epoch (2 years post-peak) is statistically significant (at the 8-level, as estimated from the rms in the VLASS “Quick Look” images, Lancel).
In addition, we found a radio source at the position of AT2019aalc in the first data release of the Rapid ASKAP Continuum Survey (RACS: racs1). The source was observed in two different fields and, consequently, at two different epochs. The observed peak flux densities of mJy/beam (on 2019-04-24) and mJy/beam (on 2020-04-30) correspond to a MHz radio luminosity of and indicate that the radio flare observed in the VLASS data started at least days after the peak of the first optical flare, sometime between April and November .
The radio luminosity in the beginning of our ATCA monitoring was at the same level ( ) when the source was observed for the second time in the VLASS in April 2021. Since the radio brightness does not show rapid variability during the months of our ATCA monitoring, we can reasonably assume that the sharp increase resulted in a long-lasting plateau. We present the ATCA radio multi-frequency light curves of AT2019aalc together with the archival observations in the middle panel of Fig. 3.
@@ 4.4.2 The radio spectrum
To further characterize the radio emission of AT2019aalc, we derived the radio spectral index of the radio source in each epoch of our ATCA monitoring. The spectral index between GHz and GHz (, defined as ) does not show any obvious signs of variability on a timescale of months. We give an average spectral index of . We calculated a pre-flare spectral index using the VLASS epoch and RACS epoch flux densities. The spectral index we derived this way is .
@@ 4.4.3 VLBI detection
The naturally-weighted VLBI map of AT2019aalc is presented in Fig. 9. We detected a single radio-emitting feature with a signal-to-noise ratio (SNR) of . We obtained the following parameters of the fitted component: a total flux density of mJy and a size of mas at a position angle of . We calculated the uncertainties of the flux density and size parameters following the formulas of fomalont. The minimum resolvable size () of a Gaussian component fitted to naturally weighted VLBI data was calculated following 2005AJ....130.2473K. This yields mas, implying that the fitted Gaussian component is resolved.
We calculated the brightness temperature of the radio-emitting feature found at the position of AT2019aalc using the following equation (Tb; Tb2):
where is the flux density of the fitted Gaussian component measured in Jy, and are the major and minor axes full width at half maximum (FWHM) of the fitted component in mas, and is the observing frequency in GHz. The obtained brightness temperature is K at GHz. The brightness temperature is an effective parameter that is commonly used in radio astronomy to describe the physical properties of emitting material in astrophysical objects (e.g., 2015A&A...574A..84L).
@@ 4.5 Spectroscopic results
We present the archival and the newly obtained optical spectra of AT2019aalc in Fig. 17, marking the notable emission lines. In the following, we highlight the remarkable line detection of the pre-flare spectrum observed before the first optical flare, the one observed between the two flares (post-flare-1 spectrum) and the five spectra taken during the second flare (over-flare-2 spectra).
@@ 4.5.1 Pre-flare spectrum
A spectrum of the pre-flare system was observed in April by the SDSS spectrograph. We present this spectrum with the best fits for the stellar and gas components in Fig. 18.
We identify strong Balmer emission lines. We calculated FWHM(Balmer) kms-1 for the broad component of the Balmer lines in the pre-flare spectrum, which is typical for an unobscured, broad-line AGN. The line widths (FWHM(Balmer) km s-1) presented in the SDSS DR7 broad-line AGN catalog (2019ApJS..243...21L) for our source are consistent with this classification.
Several forbidden lines such as [O iii] and [O iii] , [S ii] , [S ii] and [N ii] were detected. These lines are produced in the narrow-line region (NLR) in AGN. We detected the [O ii] [O ii] line doublet, which is a potential indicator of AGN-driven outflows (e.g., 2020A&A...644A..54S).
The pre-flare spectrum also indicates the detections of two high-ionization coronal lines ([Fe x] and [Fe xiv]) which are further discussed below.
@@ 4.5.2 Post-flare-1 spectrum
The first LRIS spectrum was observed almost two years after the first optical flare and two years before the second one.
@@
We identify the Bowen Fluorescence lines O iii and He ii .
The high ionization coronal lines [Fe vii], [Fe x], [Fe xiv] and [Fe vii] were detected. Furthermore, we identify the [Ne V] emission lines at and in the spectrum. Note that the pre-flare SDSS spectrum does not cover these wavelengths, therefore these lines are not necessarily related to the optical flares discussed here. These neon lines are prominent emission lines originating from the inner NLR excited by the AGN via photoionization or shocks, and are sometimes considered as coronal lines (e.g., 2020ApJ...895..147C).
@@ 4.5.3 Over-flare-2 spectra
We performed five optical spectroscopic observations of AT2019aalc during the second flare. These are characterized by a blue continuum with several more emission lines than the spectra discussed above. We compare the second LRIS spectrum of AT2019aalc (covering the broadest wavelength range) with a composite SDSS spectrum based on Seyfert 1 galaxies published in 2017MNRAS.465...95P, in Fig. 10. This comparison shows similarities but also remarkable differences discussed below.
Over the second optical flare, strong Balmer lines were detected, with two of them (H and H) appearing only in the over-flare-2 spectra.
Over the second optical flare of AT2019aalc, we identified the Bowen fluorescence line complex N iii +He ii . The complex was not detected in the earlier spectra. The relative strength of the BF complex to H reached a maximum on 2023-07-11, very close to the optical continuum peak (2023-07-13). As we detected the Bowen line N iii even days after the optical peak of the second flare, we conclude that it persists over long timescales. We also find hints of O iii, which is also known as a Bowen line (2007A&A...464..715S), however, due to a possible blending from Fe vii, its detection is questionable. Nevertheless, this Bowen line is created through the same channel (O1) as the above-discussed O iii , which supports the hint of detection. O iii was detected in some metal-rich TDEs before (2019ApJ...887..218L), such as AT2019dsg (2021MNRAS.504..792C).
We detect the high-ionization coronal lines idenfitied in the post-flare-1 LRIS spectrum over the second flare in each spectrum. Furthermore, a new coronal line appears: [Fe xi] , which has an ionization potential of (e.g., 2022ApJ...936..140R). We also detected the [Ne V] emission lines at and in the second LRIS spectrum, as well as in the post-flare-1 spectrum.
@@ 5.1.1 Balmer lines
AT2019aalc exhibits prominent Balmer lines in each spectrum. In addition, a more complex structure of these lines appears in the late-time over-flare-2 spectrum. The Balmer series lines here exhibit a redward wing which however, in some cases, might result from blending with other lines, such as H and the BF line N iii . Interestingly, the post-flare-I spectrum shows a similar structure to the Balmer line series.
@@ 5.1.2 Bowen fluorescence lines
Wavelength coincidences between emission lines (”line fluorescence”) can be an important source of radiative excitation. The Bowen Fluorescence mechanism in astrophysical environments is a good indicator of absorbed EUVsoft X-ray flux (at wavelengths shorter than the He ii Lyman limit of eV), as these high-energy photons are converted into the He ii Lyman emission required for the excitation of the O iii and N iii lines in the optical and NUV regimes (1985ApJ...299..752N; 1990agn..conf...57N). This is consistent with the multiple soft X-ray flares and the extreme UV luminosities of AT2019aalc. The BF line O iii has been observed in the Bowen Fluorescence Flares AT2017bgt (Trakhtenbrot) and AT2021loi and only in a few Seyfert galaxies before (e.g., 1986ApJ...310..679M; 1990ApJ...362...74S). However, we note that due to the atmospheric cutoff at these lines are difficult to detect. This line is typically weaker in regular Seyfert galaxies than in the case of AT2019aalc (see Fig. 10). The BF line N iii is not generally seen in AGN (e.g., 2001AJ....122..549V), however, it has been observed in several optical TDEs (spectra2) and BFFs (Trakhtenbrot; AT2021loi).
The detection of a BF line almost two years after the first optical flare of AT2019aalc suggests its long persistence, similar to the BFFs studied by Trakhtenbrot. In the case of AT2017bgt, this line was still detected days after the discovery of the transient. The Bowen features in F01004-2237 were detected even years after the initial flare (f01). We compare the over-flare-2 high-resolution LRIS spectrum of AT2019aalc to the spectra of BFFs, a TDE-Bowen, and a repeating partial TDE in Fig. 11 around the N iii He ii complex. We also compare the second LRIS spectrum of AT2019aalc with the LRIS spectrum of AT2021loi in Fig. 12, zooming into the O iii He ii region.
The He ii emission line is commonly detected in AGN and known as a recombination line, tracing ionized gas in the vicinity of the central SMBH responsible for the photoionization (e.g. 2020MNRAS.498.4550W). This line is known to be created by photoionization due to (soft) X-ray photons with a rough correspondence of He ii photon emitted for each keV X-ray photon (1986Natur.322..511P; 2019A&A...622L..10S; 2021MNRAS.504..792C). In the spectra of AT2019aalc, He ii is detected together with strong N iii and O iii BF lines and He ii , indicating its relation to the Bowen Fluorescence mechanism. Furthermore, we calculated a maximum line intensity ratio of , which significantly exceeds the typical ratios estimated for AGN (earlier 2001AJ....122..549V estimated ratios of while 2012MNRAS.421.1043S gives ). Our value is, however, more comparable with the line ratio estimated for the BFF AT2017bgt (, Trakhtenbrot).